Shift report 20011103 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 03 2001 - 15:58:33 EST

  • Next message: Apache: "Shift report 20011104 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Nov  3 15:58:33 2001, Shift: 08:00-16:00
    Supervisor: CEJ, with: BHS
    Log book#10 pages 91 - 92
    
    Shift Summary:
    ------------------------------------------------------------
    1. No usefull beam. Not-cogged beam from 12:35 to 13:15. 
       Several ramp failures.
    
    2. Run 5646 : 6k tr-6 (not-cogged beams).
       Pedestal run 5647, TPC pedestal run 5648.
    
    3. No problems reported. FV played with one of the DC 
       discriminators to investigate cross-talk between
       anode channels.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 03 2001 - 18:14:16 EST